# Rethinking Reuse in Dependency Supply Chains: Initial Analysis of NPM packages at the End of the Chain

### v2

@@ Abstract.
The success of modern software development can be largely attributed to the concept of code reuse, such as the ability to reuse existing functionality via third-party package dependencies, evident within massive package networks like NPM, PyPI and Maven. For a long time, the dominant philosophy has been to ‘reuse as much as possible, without thought for what is being depended upon’, resulting in the formation of large dependency supply chains that spread throughout entire software ecosystems. Such heavy reliance on third-party packages has eventually brought forward resilience and maintenance concerns, such as security attacks and outdated dependencies.
In this vision paper, we investigate packages that challenge the typical concepts of reuse–that is, packages with no dependencies themselves that bear the responsibility of being at the end of the dependency supply chain. By avoiding dependencies, these packages at the end of the chain may also avoid the associated maintenance and resilience risks typically beyond their control. For our initial analysis, we look at a sample of heavily-depended upon NPM packages, and identify that such end-of-chain packages make up a significant portion of these critical dependency chain (over 50%). We find that these end-of-chain packages vary in characteristics and not just packages that can be easily replaced: an active, well-maintained package at the end of the chain; a ”classical” package that has remained unchanged for 11 years; a trivial package nested deep in the dependency chain; a package that may appear trivial; and a package that bundled up and absorbed its dependencies.
We then ask ourselves: Should project maintainers follow suit and become less dependent on third-party packages? What would the impact on supply chains be if any of these critical packages were abandoned or attacked by bad actors? The vision of this paper is to advocate for a shift in software development practices toward minimizing reliance on third-party packages, particularly those at the end of dependency supply chains. We argue that these end-of-chain packages offer unique insights, as they play a key role in the ecosystem. At the same time, we do not want to erase the success of NPM reuse, but we need to understand how to strategically reuse so that the benefits are fully realized by the maintainer and the ecosystem as a whole.
@@ 1. Introduction
The practice of reusing code via third-party packages is a common part of software development, evident by the popularity of package managers such as Node Package Manager (NPM), hosting over 2.1 million packages (or packages) as of September 2022 (The OpenJS Foundation, [n. d.]). The abundance of such packages enables developers to integrate already-made solutions over writing their own implementations, with well-established benefits for productivity, code quality and defect reduction (Lim, 1994; Basili et al., 1996).
Despite the ubiquity of such code reuse, there is not a lack of risk; recent work has focused on the impact of dependencies on the security and sustainability of software projects (Decan et al., 2018; Wattanakriengkrai et al., 2022), especially in regards to resilience and maintenance. Many NPM packages depend directly on a number of other packages, that then have their own dependencies and so on, resulting in lengthy chains of indirect dependencies. Prior work has considered code reuse as a double-edged sword (Gkortzis et al., 2019): the maturity and open-source nature of these packages can make systems more secure, but at the same time the increased attack surface can introduce vulnerabilities. Maintainers and users of a package may not necessarily be aware of the full scale and impact of a package’s dependency chain - the result is opportunities for unsecured links on the supply chain to introduce security and maintenance issues to upstream dependent packages.
The unpublishing of the left-pad package in 2016 subsequently broke numerous other important packages on NPM, including software maintained by large companies such as Meta (Collins, 2016). While Meta’s React package didn’t directly depend on this trivial, 11 line package, it was impacted via its dependency chain. Increasingly, there are also social aspects such as protestware to consider - the introduction of potentially malicious code for political or financial reasons (Kula and Treude, 2022). In the case of the node-ipc package, when ran by users from Russia or Belarus, the code would attempt to replace all files with a heart emoji in protest of the war in Ukraine. In response to these concerns and high-profile cases that propagate across an entire ecosystem such as Maven’s Log4Shell vulnerability (YfryTchsGD, 2021), efforts to tighten the software supply chain have emerged. The Alpha and Omega Project specifically aims to help the most critical open source projects improve their security postures (Foundation, 2024). More and more, software developers are looking to reduce dependencies and their overhead (Lu, 2024).
@@
Despite the popular culture of reuse that is apparent within the NPM ecosystem, some packages curiously abstain from the use of third-party dependencies themselves, making them the end of the dependency supply chain. In doing so, these packages may also be avoiding the risk factors that come with dependency use. Previous work investigating the NPM ecosystem has focused on the impact of lengthy dependency chains, but not necessarily the reverse, and the results of these studies may not generalise to the no dependency case. In this vein, there may be undiscovered benefits to refraining from code reuse in specific scenarios wherein the ‘buck’ of the supply chain stops with this package.
@@
In this paper, we argue that analysis of software ecosystems like NPM should not ignore the role of end-of-chain packages, and how the unforeseen benefit may stack up against the productivity benefits of code reuse. For the most popular packages on NPM, our preliminary results indicate that packages with no dependencies are indeed evident in the dependency chain, comprising just over 50% of the dependency chain in our sample. We select a series of motivating cases, which we believe characterise some the different types of end-of-chain dependency packages that exist in the NPM ecosystem: popular packages, ‘classical’ packages, trivial packages, complex trivial packages and packages that absorb their dependencies. From these cases, we present a series of open research questions to guide researchers in investigating, characterising and understanding the perils and best practice of these no dependency packages, their impact and the motivations behind reducing dependencies. Our new idea is to explore how understanding the packages at the end of the chain may impact the quality and resilience of the supply chain itself, and further reveal more insights into the complicated trade-offs of code reuse in software development.
2. Extracting End of Chain from the Top 10 NPM packages
@@
We conduct an initial analysis of the top ten most depended upon NPM packages. We begin by looking at the top 10 most depended upon packages in the NPM ecosystem and their chain of upstream dependencies. Being one of the largest and most studied software ecosystems (The OpenJS Foundation, [n. d.]; Zimmermann et al., 2019; Kula et al., 2017), we choose to limit this initial analysis to only the NPM ecosystem. We select this subset of packages because they represent a sample of the packages critical to the NPM dependency chain. First, we utilise the packages.io [^1] https://packages.io/ API to get the top 10 packages sorted by dependency count. Next, for those 10 packages, we mine the NPM registry [^2] https://registry.npmjs.org for dependency data, and recursively build their chain of dependencies; that is, the packages they depend upon. We look only at the dependencies field on the NPM registry, and ignore devDependencies and peerDependencies. devDependencies pertain to only dependencies required during development of the plug-in which are not installed in production, while peerDependencies often denotes that a package is a plug-in for another package, but it may or may not actually use APIs from that package. We limit our analysis to regular dependencies to avoid the complexity of these cases for this investigation. Additionally, no recursive dependencies are present in this dataset. Table 1 shows the final dataset statistics.
@@ 2.1. Prevalence of end-of-chain packages
3. Exploring Five Cases of End-of-Chain
@@
Based on the dataset results, we now present five motivating cases of no dependency packages at the end of the dependency chain, their impacts on the dependency chain and the implications of their lack of dependencies. We select these cases from our sample based on their prospective interest to the software engineering community. The first case describes the case of an active, well-maintained package at the end of the dependency chain. The second case describes the case of a ‘classical’ package that has remained unchanged for 11 years, that’s lack of activity may be a side effect of its simplicity and lack of dependencies. In the third case, we take a look at the case of a trivial package nested deep in the dependency chain of the popular eslint package. Forth, we observe the case of a package that may appear trivial, but package history and activity suggests otherwise. Finally, we look at the case of a package that bundled up and absorbed its dependencies.
3.1. Case 1: The popular package (Most Depended Upon)
@@
The most depended upon NPM package according to packages.io data also has zero dependencies: the typescript package [^3] https://github.com/microsoft/TypeScript. As many projects utilize TypeScript, the popularity of this package is not surprising. The package is often installed as part of the build process for a package, where TypeScript will be transpiled into JavaScript. The package is 32MB and the GitHub repository is extremely active, with a commit in the last two days, 5000+ issues and 338 pull requests. The repository also lists 761 contributors. There is also a level of expected support, as the package is maintained by Microsoft. In terms of security, the repository has a security policy and no reported vulnerabilities on Snyk (Snyk, 2024b). A look at the history of this package’s dependencies reveals that it has always had zero dependencies.
We look at this case as an example of an active, popular package at the end of the dependency chain. Due to its position in the JavaScript ecosystem, a lack of dependencies may be a conscious decision to secure the supply chain. Motivations like these are the kind we aim to reveal with further work.
3.2. Case 2: The ‘classical’ package (11yr last updated)
@@
Some packages don’t necessitate updates and unlike the previous example, may appear inactive; we call these ‘classical’ packages. The package imurmurhash [^4] https://github.com/jensyt/imurmurhash-js is an implementation of the MurmurHash3 hashing algorithm, last updated 11 years ago, with a single maintainer. Despite these factors, it still has 23.6 million users. There is only a single, closed issue concerning a false positive virus detection (dalerka, 2020); however, the maintainer still appears responsive.
One user in the issue suggests that based on the age of the package, and that the virus detection issue continues to exist, people should no longer use it in favour of more active packages.
@@
This case exemplifies the ‘classical’ package case, as the implementation of an algorithm to a specification is unlikely to require the kinds of updates other packages do, especially when there are no dependencies to update. We use this example to illustrate that by reducing dependencies, developers may be able to reduce required maintenance.
3.3. Case 3: The trivial package (is-number)
@@
Whether the optimisation changes would be impactful or not, this package is clearly a case of the dependency chain relying on a trivial, inactive package with only a few contributors. Unlike the previous case of a classical package, however, there are possibly needed updates. The observation from is-number is that not every package with no dependencies is useful; in fact, is-number may be an obvious candidate for removal in projects looking to reduce dependencies themselves.
3.4. Case 4: Complex trivial package (internally vulnerable)
@@
This package is an example of how small, seemingly simple packages can introduce weak points in the dependency chain. In this case, the problems with this package stem from the complexities of regular expressions and string parsing. Due to the single functionality, lack of dependencies and small size, this package may appear deceptively trivial to potential users. The use of these types of packages may not necessarily be unwarranted though; by reusing a package, vulnerabilities can be reported, consumers of a package can be notified, and updates can be pushed.
3.5. Case 5: The bundled package (clone-and-own)
We present another interesting case, where a zero dependency package absorbed its dependencies. The chalk package [^7] https://github.com/chalk/chalk, for styling terminal text using ANSI, has no dependencies, but this was not always the case. Prior to removing all dependencies, the package had two dependencies: supports-color and ansi-styles, both also maintained under the GitHub chalk profile. In 2021, the packages were bundled; their files were moved into the vendor folder of the chalk package (sindresorhus, 2021), while the dependencies are still maintained separately. In the GitHub issue discussing the roadmap for this update, the main contributor explains their reasoning (sindresorhus, 2018):
@@ 3.5. Case 5: The bundled package (clone-and-own)
While the developer decided not to go through with bundling at the time, they later did so anyway, though further discussion cannot be found on the repository. Cases like these may help us understand how and when developers choose to reduce dependencies.

### v3

# Towards Sustainable and Secure Reuse in Dependency Supply Chains: Initial Analysis of NPM packages at the End of the Chain

@@ Abstract.
Much of the success of modern software development can be attributed to code reuse. The ability to reuse existing functionality via third-party dependencies has enabled massive gains in productivity, but for a long time the dominant philosophy has been to ‘reuse as much as possible, without thought for what is being depended upon’, creating fragile dependency chains. Heavy reliance has raised resiliency and maintenance concerns. In this vision paper, we investigate packages that challenge the typical concepts of reuse–that is, packages with no dependencies themselves that bear the responsibility of being at the end of the dependency supply chain. By avoiding dependencies, these packages at the end of the chain may also avoid the associated risks. Our initial analysis of the most depended upon NPM packages shows that such end-of-chain packages make up a significant portion of these critical dependency chain (over 50%). We find that these end-of-chain packages vary in characteristics and are not just packages that can be easily replaced, and present five cases. We then ask ourselves: Should maintainers minimize external dependencies? We argue that these packages reveal important lessons for strategic reuse—balancing the undeniable benefits of dependency ecosystems with sustainable, secure practices.
Software Engineering, Software Ecosystems, Security and Maintenance
conference: International Workshop on Software Engineering in 2030; a; accs: Software and its engineering Software libraries and repositoriescopyright: acmcopyright
@@ 1. Introduction
Douglas McIlroy proposed the idea of mass-produced, component based software engineering in 1968 (McIlroy et al., 1968). He envisioned a future where software would be assembled from reusable parts rather than repeatedly built from scratch, akin to mass manufacturing. Today, over 50 years later, the mindset of ‘use over build’ dominates the industry; almost no software is built entirely from the ground up, instead making use of vast collections of already-existing software libraries. The Node Package Manager (NPM), one of the largest ecosystems of software libraries, hosts over 3.5 million NPM ‘packages’ as of July 2025 (NPM, [n. d.]), each with their own complicated chains of dependency. Though the benefits of library use for productivity, code quality and defect reduction are well established (Lim, 1994; Basili et al., 1996), recent focus has shifted to the security and sustainability of software projects (Decan et al., 2018; Wattanakriengkrai et al., 2022), especially in regards to resilience and maintenance. Prior work has considered code reuse as a double-edged sword (Gkortzis et al., 2019): the maturity and open-source nature of these packages can make systems more secure, but at the same time the increased attack surface can introduce vulnerabilities. McIlroy’s metaphor of industrialization, suppliers and supply chains remains just as relevant to software today as when it was proposed.
Recently, numerous high-profile cases of weaknesses in the software supply chain have gained attention from the software engineering community due to their widespread impact. The Java Log4Shell vulnerability (YfryTchsGD, 2021) impacted millions of users, with 60% of affected libraries vulnerable via indirect dependencies (Tal, 2021). In 2016, the unpublishing of the single function left-pad package lead to NPM changing policies on library deletion, after large tech companies such as Meta and PayPal were left suddenly unable to build projects. Increasingly, there are also social aspects such as protestware to consider - the introduction of potentially malicious code for political or financial reasons (Kula and Treude, 2022). These cases highlight the challenge posed by lengthy and complex chains of indirect dependencies that have emerged across software ecosystems such as NPM. Maintainers and users of a package may not necessarily be aware of the full scale and impact of a package’s dependency chain - the result is opportunities for unsecured links on the supply chain to introduce security and maintenance issues to upstream dependent packages. Additionally, they may not be aware of who maintains the software they rely on - often open source software is maintained by small teams of volunteers. Like sourcing parts from unknown factories, the promise of mass-produced, reliable components has not been realized.
Figure 1. Example of a dependency chain for the NPM package axios [^1] https://www.npmjs.com/package/axios. Darker nodes indicate direct dependencies, lighter nodes indicate indirect dependencies, and crosses indicate end-of-chain packages.
@@
Standing against these trends are projects that abstain from the use of third-party dependencies entirely. In doing so, these packages may also be avoiding the risk factors that come with dependency use. Previous work investigating the NPM ecosystem has focused on the impact of lengthy dependency chains, but not necessarily the reverse, and the results of these studies may not generalise to the no dependency case. In this vein, there may be undiscovered benefits to refraining from code reuse in specific scenarios wherein the ‘buck’ of the supply chain stops with these packages.
In this paper, we preform an initial analysis of the ten most depended upon libraries in the NPM ecosystem and their dependency chains (consisting of 216 unique libraries), in order to identify and characterise the types of dependencies that exist at the end of the chain. For the most popular packages on NPM, our preliminary results indicate that packages with no dependencies are indeed evident in the dependency chain, comprising just over 50% of the dependency chain in our sample. We select a series of motivating cases, which we believe characterise some the different types of end-of-chain dependency packages that exist in the NPM ecosystem: 1) core ecosystem packages that are critical to the ecosystem; 2) ‘classical’ packages that don’t necessitate updates; 3) unmaintained trivial packages; 4) ‘not trivial‘ trivial packages that appear deceptively simple; 5) and bundled packages that absorb their dependencies. We believe these cases require further study and are of interest to the software engineering community.
For a long time, the dominant philosophy has been to ‘reuse as much as possible, without thought for what is being depended upon’. We propose an alternative: sustainable, secure reuse. In line with this, we pose the following overarching question:
How can we strive towards more sustainable reuse, and can we learn from libraries that lack dependencies?
@@
More and more, software developers are looking to reduce dependencies and their overhead (Lu, 2024). The Alpha and Omega Project specifically aims to help the most critical open source projects improve their security postures (Foundation, 2024). In this paper, we argue that analysis of software ecosystems like NPM should not ignore the role of end-of-chain packages, and how the unforeseen benefit may stack up against the productivity benefits of code reuse. We present a series of open research questions to guide researchers in investigating, characterising and understanding the perils and best practice of these no dependency packages, their impact and the motivations behind reducing dependencies, within the scope of the following topics:
  • Characterising End-Of-Chain Packages: These questions aim to better understand the structural traits and behaviour of packages with no dependencies, including their size, update frequency, and role in the ecosystem.
  • Understanding Developer Motivations: These questions explore why developers create or adopt zero-dependency packages, such as the security, performance and maintenance factors that might be behind their choices.
  • Evaluating the Risks, Impact and Benefits of No Dependencies: These questions assess the reuse trade-offs of reducing dependencies on security, maintenance effort and ecosystem stability.
  • Rethinking Reuse for the Ecosystem: These questions reframe reuse strategies to prioritize sustainability and resilience, with implications for governance, adoption, and long-term ecosystem health.
Overall, we envision a future where software library reuse is secure, sustainable and intentional, and where developers understand the impact and scope of their supply chains. In order to strive towards this goal, however, we must understand and study reuse from the perspective of libraries that minimise reuse. Additionally, we want to understand the role of these end-of-chain libraries on the quality and resilience of the supply chain itself, and further reveal more insights into the complicated trade-offs of code reuse in software development.
2. How prevalent are end-of-chain libraries for the Top 10 NPM Packages?
@@
In order to understand the nature of end of chain libraries, we conduct an initial analysis of the top ten most depended upon NPM libraries and their dependency chains. Being one of the largest and most studied software ecosystems (NPM, [n. d.]; Zimmermann et al., 2019; Kula et al., 2017), we choose to limit this initial analysis to only the NPM ecosystem. We select this subset of packages because they represent a sample of the packages critical to the NPM dependency chain. First, we utilise the libraries.io [^2] https://libraries.io/ API to collect the top 10 packages sorted by dependency count. Next, for those 10 packages, we mine the NPM registry [^3] https://registry.npmjs.org for dependency data, and recursively build their chain of dependencies; that is, the packages they depend upon.
For this initial analysis we look only at the dependencies field on the NPM registry, and ignore devDependencies and peerDependencies. devDependencies pertain to only dependencies required during development of the library that are not installed in production or by dependent libraries. These typically include testing frameworks and programming aids like linters. APIs from these dependencies are not called in any production code. Similarly, peerDependencies often denotes that a package is a plug-in for another package, but it may or may not actually use APIs from that package. We limit our analysis to regular dependencies to avoid the complexity of these cases for this investigation. Additionally, no recursive dependencies are present in this dataset. Table 1 shows the final dataset statistics.
@@
Libraries with no dependencies make up a majority of the supply chain for the most depended upon NPM packages (52.78%), however, they remain understudied.
3. Motivating Examples: Exploring Five Cases of End-of-Chain Libraries
From the dataset we select five motivating cases of libraries with no dependencies that exist at the end of the dependency chain, their impacts and the implications of their lack of dependencies. We select these cases from our sample based on their prospective interest to the software engineering community.
@@ 3. Motivating Examples: Exploring Five Cases of End-of-Chain Libraries
The first case describes the case of an active, well-maintained package at the end of the dependency chain. The second case describes the case of a ‘classical’ package that has remained unchanged for 11 years, that’s lack of activity may be a side effect of its simplicity and lack of dependencies. In the third case, we take a look at the case of a trivial package nested deep in the dependency chain of the popular eslint package. Forth, we observe the case of a package that may appear trivial, but package history and activity suggests otherwise. Finally, we look at the case of a package that bundled up and absorbed its dependencies.
3.1. Case 1: The Core Ecosystem Package
The most depended upon NPM package according to libraries.io data also has zero dependencies: the typescript package [^4] https://github.com/microsoft/TypeScript. Given the popularity of the TypeScript language, this isn’t surprising. The package is often installed as part of the build process for a package, where TypeScript will be transpiled into JavaScript. As of July 2025, 1,063,224 of the 3,583,585 libraries on the NPM registry make use of the typescript library - almost 1/3 of the entire ecosystem. This makes typescript a critical part of infrastructure in the NPM ecosystem.
@@ 3.1. Case 1: The Core Ecosystem Package
The GitHub repository for the typescript package is extremely active, with recent commits, 5000+ issues, 338 pull requests and a large list of 761 contributors. As the package is maintained by Microsoft, there is also an expected level of support. In terms of security, the repository has a security policy and no reported vulnerabilities on Snyk (Snyk, 2024b). A look at the history of this package’s dependencies reveals that it has always had zero dependencies. Comparatively, the library’s unpacked size is relatively large at 32MB (when most packages are measured in KB).
Some libraries that play an important role in the NPM ecosystem, such as typescript, have no dependencies. Due to their position in the Node.js ecosystem, such core ecosystem libraries may be more likely to have smaller supply chain footprints as a conscious decision to reduce supply chain risk and maintain more control. Additionally, it may be that libraries that remain self-contained are more likely to see widespread adoption by software developers. Additionally, it’s size and complexity show that even large projects can avoid dependencies. These observations raise important questions about the motivations and characteristics of end-of-chain libraries. Core Ecosystem Packages are likely to be characterised by active maintenance, high dependent count and large communities.
3.2. Case 2: The ‘Classical’ Package
@@
The imurmurhash [^5] https://github.com/jensyt/imurmurhash-js package is an interesting case; an implementation of the MurmurHash3 hashing algorithm, last updated 11 years ago, with a single maintainer. Despite these factors, it still has 23.6 million users and an active maintainer. Due to the nature of the library - an implementation of an algorithm to a specification - there is little need for the kinds of updates other libraries require, especially with no dependencies to update. In some sense, these libraries may be ’classical’.
There is only a single, closed issue concerning a false positive virus detection (dalerka, 2020).
@@
Some libraries may decrease the need for maintenance by avoiding dependencies. Additionally, there may be cases of libraries that are straightforward to implement without dependencies. This case raises interesting questions concerning the risks and benefits of using no dependencies, as well as insights about what kinds of libraries can avoid dependencies. Classical Packages are likely to be characterised by lack of recent updates, but maintainers that respond to issues and no active vulnerability reports for the latest version.
3.3. Case 3: The Unmaintained Trivial Package
@@
Whether the optimisation changes would be impactful or not, this package is clearly a case of the dependency chain relying on a trivial, inactive package with only a few contributors. Unlike the previous case of a classical package, however, there are possibly needed updates.
Not every library at the end of the chain is useful; in fact, unmaintained trivial packages like is-number may be an obvious candidate for removal in projects looking to reduce dependencies themselves. Libraries such as is-number also raise questions about the types of libraries that can avoid dependency use. Unmaintained trivial packages are likely to be characterised as trivial packages with no recent commits, inactive maintainers and stale issues and/or pull requests.
3.4. Case 4: The ’Not Trivial’ Trivial Package
@@
This package is an example of how small, seemingly simple packages can introduce weak points in the dependency chain. Due to the single functionality, lack of dependencies and small size, this package may appear deceptively trivial to potential users. The use of these types of packages may not necessarily be unwarranted though; by reusing a package, vulnerabilities can be reported, consumers of a package can be notified, and updates can be pushed. This package shows the risks and benefits of code reuse, in regards to vulnerabilities, but also community reporting of those vulnerabilities. ‘Not trivial’ trivial packages are likely to be characterised as trivial packages with active maintenance, active community (issues and PRs), history of vulnerabilities and high contributor counts.
3.5. Case 5: The Bundled Package
We present another interesting case, where a zero dependency package absorbed its dependencies. The chalk package [^8] https://github.com/chalk/chalk, for styling terminal text using ANSI, has no dependencies, but this was not always the case. Prior to removing all dependencies, the package had two dependencies: supports-color and ansi-styles, both also maintained under the GitHub chalk profile. In 2021, the packages were bundled; their files were moved into the ‘vendor’ folder of the chalk package (sindresorhus, 2021), while the dependencies are still maintained separately. In the GitHub issue discussing the roadmap for this update, the main contributor explains their reasoning (sindresorhus, 2018):
@@ 3.5. Case 5: The Bundled Package
While the developer decided not to go through with bundling at the time, they later did so anyway, though further discussion cannot be found on the repository.
Cases like these may help us understand how and when developers choose to reduce dependencies, and specifically, when they choose to bundle over other possibilities. The frequency of this technique is currently unknown, as are its risks and benefits. Bundled packages are likely to be characterised as packages with a previous history of dependencies, with dedicated directories such as ‘vendor’.
4. Discussion
Our results suggest that end-of-chain packages potentially play an important role in the NPM ecosystem, with four of the ten most depended upon packages on the dependency chain having no dependencies themselves. As summarized in Table 2, we identify five categories of end-of-chain libraries, their importance to the software ecosystem and researchers, and potential indicators. While there are obvious benefits to a lack of dependence on other software, such as a lack of inherited vulnerabilities and smaller project footprint, other benefits remain unknown. Additionally, we don’t know how these gains stack up against the productivity benefits of code reuse. By looking at five example cases of packages with no dependencies at the end of the supply chain, we make the case that there may be new insights to be gained from similar packages across software ecosystems. Additionally, End-of-chain packages may have their own risk that need to considered, and the motivations behind the use of no dependencies are currently unstudied. We present a research agenda for understanding these aspects of end-of-chain packages.
Table 2. Potential taxonomy of the end-of-chain libraries
Category Importance Possible Characteristics
Core Ecosystem Packages Packages that are critical to the ecosystem. Active Maintenance
High Dependent Count
Large Community
Classical Packages Packages that remain stable, but may be misidentified as inactive. No Recent Commits
Responsive Maintainer
No Active Vulnerability Reports
Unmaintained Trivial Packages Unmaintained trivial packages that might be candidates for removal. Small Size
No Recent Commits
No Maintainer Activity
Open Issues and PRs
Not Trivial Trivial Packages Packages that solve deceptively simple tasks and require care. Small Size
Active Maintenance
Issue Activity
History of Vulnerabilities
High Contributor Count
Bundled Packages Packages that chose to remove dependencies and instead bundle. History of Dependencies
Existence of ‘vendor’ Directory.